Find the equation of a line with a slope of 4 and passing through(-2,7)
Answer by Alan3354(69443)

Find the equation of a line with a slope of 4 and passing through(-2,7)
0 solutions
--------------------
y-y1 = m*(x-x1) where (x1,y1) is the point and m is the slope
y-7 = 4*(x+2)
y-7 = 4x+8
y = 4x+15 Slope-intercept form
4x-y = -15 standard form